OpenStudy (anonymous):

3x+2y-3z=9 2x-3y+2z=15 4x-2y+4z=18

OpenStudy (anonymous):

x+y+2z=18 x+y+4z=28 x+9y-3z=33

OpenStudy (anonymous):

6x-y+3z=-15 x+3y-z=4 3x+3y-4z=-8

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

a broad way theater has 600 seats divided into orchestra, main and balcony seating. Ocrcheastra seats sell for $50, main sears for $40 and balcony seats for $25. If all the seats are sold the gross revenue to the theater is $22,200. If a ll the main and balcony seats are sold but only half the orchestra seats are sold. the gross revenue is $19,200. How many are there of each kind of seat.

OpenStudy (anonymous):

for the second question i got z=5

OpenStudy (anonymous):

2nd z = 5 is correct

OpenStudy (anonymous):

i got the solution for the 1st 3, i will go get a cup of it before starting on the 4th

OpenStudy (anonymous):

already?

OpenStudy (anonymous):

hehe they are piece of cake

OpenStudy (anonymous):

im starting the 4th

OpenStudy (anonymous):

yeah my daughter is bugging me and the tv is loud.

OpenStudy (anonymous):

im stucks/.

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

the second one.

OpenStudy (anonymous):

-8y-12z=-129

OpenStudy (anonymous):

oh thats suppose to be -21z

OpenStudy (anonymous):

-8y-21z=17

OpenStudy (anonymous):

x+y+2z=18 x+y+4z=28 ------------subtract -2z = -10 ===>z = 5 x+y+4z=28 x+9y-3z=33 ------------subtract -8y +7z = -5 -8y + 5*7 = -5 -8y = -5 - 35 -8y = -40 y = 5

OpenStudy (anonymous):

now you got z and y, what is x?

OpenStudy (anonymous):

you passed your 12mins a question

OpenStudy (anonymous):

i know kids are annoying

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

x+y+2z=18 x + 5 + 2*5 = 18 x + 15 = 18 what is x

OpenStudy (anonymous):

3

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so its 3,5,5 for the second question correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6x-y+3z=-15 x+3y-z=4 3x+3y-4z=-8

OpenStudy (anonymous):

if you tell me all the answers i will never ask anything from you again :D

OpenStudy (anonymous):

lol you little cheater

OpenStudy (anonymous):

eh..

OpenStudy (anonymous):

i wont so u can ask more questions :P

OpenStudy (anonymous):

i still have a two hour test to take before 11

OpenStudy (anonymous):

x =-3 y = 3 z = 2

OpenStudy (anonymous):

i got the answer for the 4th one

OpenStudy (anonymous):

the word problem im working on that.

OpenStudy (anonymous):

x=120

OpenStudy (anonymous):

there is no x; its orchestra :P

OpenStudy (anonymous):

so i used O

OpenStudy (anonymous):

O = 120 M = 280 B = 200

OpenStudy (anonymous):

now you have time to do the other 2 problems

OpenStudy (anonymous):

thank you .

OpenStudy (anonymous):

x+y+z=12 2x+4y+5z=35 4x-y=24

OpenStudy (anonymous):

i have three more left

OpenStudy (anonymous):

didnt we do 3 already

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

x+y+z=12 2x+4y+5z=35 4x-y=24 y = 4; what is x?

OpenStudy (anonymous):

x =-3 y = 3 z = 2 which one is this one

OpenStudy (anonymous):

let me look; u make me do the work then you lose the solution

OpenStudy (anonymous):

lol i just saw it.

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

6x-y+3z=-15 x+3y-z=4 3x+3y-4z=-8 x =-3 y = 3 z = 2

OpenStudy (anonymous):

x+y+2z=18 x+y+4z=28 x+9y-3z=33 x = 3 y = 5 z = 5

OpenStudy (anonymous):

a broad way theater has 600 seats divided into orchestra, main and balcony seating. Ocrcheastra seats sell for $50, main sears for $40 and balcony seats for $25. If all the seats are sold the gross revenue to the theater is $22,200. If a ll the main and balcony seats are sold but only half the orchestra seats are sold. the gross revenue is $19,200. How many are there of each kind of seat. O = 120 M = 280 B = 200

OpenStudy (anonymous):

whats 3x+2y-3z=9

OpenStudy (anonymous):

3x+2y-3z=9 2x-3y+2z=15 4x-2y+4z=18 -------------- your turn. you have time :P

OpenStudy (anonymous):

x+y+z=12 2x+4y+5z=35 4x-y=24 x = 7 y = 4 z = ?

OpenStudy (anonymous):

x+y+z=12 2x+4y+5z=35 4x-y=24 x=7 y=4 z=1

OpenStudy (anonymous):

correct, now u do the last 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 7z+z=27

OpenStudy (anonymous):

what did you do

OpenStudy (anonymous):

3x+2y=3z=9 4x-2y+4z=18

OpenStudy (anonymous):

now i have 5 2x-3z=-12

OpenStudy (anonymous):

3x+2y-3z=9 4x-2y+4z=18 --------------------- 7x + z = 27

OpenStudy (anonymous):

ok 2x-3z=-12 7z+z-27

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

2x-3y+2z=15 *2 4x-2y+4z=18*3 4x - 6y + 4z = 30 12x - 6y + 12z = 54 ----------------------subtract -6x - 6z = - 24

OpenStudy (anonymous):

z=6

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

7x + z = 27 -6x - 6z = -24

OpenStudy (anonymous):

how many min u have

OpenStudy (anonymous):

10 i think

OpenStudy (anonymous):

ok u have time

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

7x + z = 27 -6x - 6z = -24 ----------------i trust u

OpenStudy (anonymous):

42x+6z=162

OpenStudy (anonymous):

42x+6z=162 -6x - 6z = -24 --------------scared u huh

OpenStudy (anonymous):

so i got 36x=138

OpenStudy (anonymous):

but it come out 3.8

OpenStudy (anonymous):

we did something wrong let me check

OpenStudy (anonymous):

how many minutes?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

6 min

OpenStudy (anonymous):

hurry

OpenStudy (anonymous):

x = 4 y = -3 z = ?

OpenStudy (anonymous):

are we doing 6x-y+3z=-15

OpenStudy (anonymous):

pellet times up we got a 60

OpenStudy (anonymous):

we did that one 6x-y+3z=-15 x+3y-z=4 3x+3y-4z=-8 x =-3 y = 3 z = 2

OpenStudy (anonymous):

i put the wrong answer in the wrong question

OpenStudy (anonymous):

3x+2y-3z=9 2x-3y+2z=15 4x-2y+4z=18 x = 4 y = -3 z = ?

OpenStudy (anonymous):

u can fix that right

OpenStudy (anonymous):

times up. it submitted it. I have totake it again.

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

what did u get?